OpenStudy (anonymous):

Hey i have 1 question... In a quadratic equation, when you solve by factoring, does it mater where you put the factoring numbers? EX: (x+5)(x-4) does it matter if i put (x+4)(x-5)

OpenStudy (jamesj):

Very definitely, as the two are not equal. (x+5)(x-4) = x^2 + x - 20 but (x+4)(x-5) = x^2 - x - 20 which are different.

OpenStudy (anonymous):

yes (x+4)(x-5) = x^2 - x - 20 and (x-4)(x+5) = x^2 + x -20

OpenStudy (anonymous):

Oh thanks alot